OpenStudy (anonymous):

Find the limits at infinity:

OpenStudy (anonymous):

\[\lim x > infinity \frac{ x }{ x{3} + 1 }\]

OpenStudy (raffle_snaffle):

CAN YOU APPLY l'HOPSITALS RULE?

OpenStudy (anonymous):

\[2. \lim x > -infinity \frac{ \sqrt{x} }{ \sqrt{x} +4 }\]

OpenStudy (raffle_snaffle):

I believe the answer to the first question is 1/3

OpenStudy (anonymous):

Is that x/(x^3+1) OR x/(3x+1) ???

OpenStudy (anonymous):

@raffle_snaffle uhm. I think there's no need because there's this rule that says that if r>0 and r is an element of Q, " lim x > infinity 1/x^r = 0 "

OpenStudy (raffle_snaffle):